Zinc metal reacts with hydrochloric acid according to the following balanced equation: Zn(s)+2HCl(aq)→ZnCl2(aq)+H2(g) When 0.119 g of Zn(s) is combined with enough HCl to make 51.3 mL of solution in a coffee-cup calorimeter, all of the zinc reacts, raising the temperature of the solution from 22.5 ∘C to 24.1 ∘C. Part A Find ΔHrxn for this reaction as written. (Use 1.00 g/mL for the density of the solution and 4.18 J/(g⋅∘C) as the specific heat capacity.) Express your answer in kilojoules per mole to two significant figures.


1
Expert's answer
2021-06-10T03:35:21-0400

Mass of solution = volume x density


= 54.5 x 1.0 = 54.5 g


Heat released by reaction = heat absorbed by solution


= mass x specific heat x temperature change of solution


= 54.5 x 4.18 x (24.7 - 22.4)


= 523.963 J


Moles of Zn = mass/molar mass of Zn


= 0.104/65.41 = 0.001590 mol


ΔH = -heat released by reaction/moles of Zn


= -523.963/0.001590


= -3.30 x 105 J/mol = -330 kJ/mol